OpenStudy (anonymous):

Expand the logarithm: log((s√7)/t^2) I know the answer to this question is log(s)+1/2log7-2log(t) but i'd like to understand WHY that is the answer, not just the steps but also how to solve this kind of problem in the future

OpenStudy (amistre64):

application of log rules log(ab) = log(a) + log(b) log (x^n) = n log(x) why?? x^n = xxxx....x, n times log(xxx...x) = log(x) + log(x) + log(x) + ... +log(x) ; n times

OpenStudy (amistre64):

a/b = ab^(-1) log(a/b) = log(ab^(-1)) = log(a) + log(b^(-1)) = log(a) - log(b)

OpenStudy (amistre64):

log(1) = log(a/a) = log(a)-log(a) = 0